Use the reduced 2nd virial coefficient to calculate molar volume of a gas.
Answer #1

Z = (PV/RT) = 1 (B/V)

This is called reduced second virial coefficient.

whereas,

z= compressibility factor

B= second virial coefficient (m3/mol)

P= pressure (Pa)

V= molar volume (m3/mol)

T= Temperature (K)

R= gas constant (8.314 j/mol k) in SI unit ->(fixed value)

For the given value of B, Pressure and Temperature we can calculate molar volume (v) of gas by rewriting the following equation.

V= RT/P *(1+B/V). In this equation, the only unknown is V. So we calculate molar volume of a gas from this equation.